What is the specific heat capacity of a 10, k, g,


10kg object that needs 2, point, 1, times, 10, to the power 5 , J,


2.1×105

J to have its temperature raised by 5, point, 0, degrees, C,


5.0∘

C?


1
Expert's answer
2021-11-17T18:31:13-0500

The specific heat capacity:


"c=\frac Q{m\Delta t}=\frac{2.1·10^5}{10·5}=4200\text{ J/(kg}·\text{K)}."